OpenStudy (anonymous):

Find dy/du, du/dx, and dy/dx y= u^2 , u=4x+7

OpenStudy (anonymous):

2u,4,8u in order as give. :)

OpenStudy (anonymous):

\[\frac{dy}{du}=2u\]\[\frac{du}{dx}=4\]\[\frac{dy}{dx}=\frac{dy}{du}\times \frac{du}{dx}=4\times2u=8u\]

OpenStudy (anonymous):

can u guys show the working because i dont understand how u got that did u use the quotient rule to find dy/du

OpenStudy (anonymous):

purple did you get and understand the solution ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Just use the normal rule for differentiating polynomials to find dy/du and du/dx and the use the chain rule to find dy/dx.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok i cant do it for you dy/dx =2u ,du/dx==4 and dy/dx= (dy/du)(du/dx)=2u(4)= 8u

OpenStudy (anonymous):

wouldnt dy/dx be 8(4x+7)

OpenStudy (anonymous):

i mean i can do it for u..lol..

OpenStudy (anonymous):

using the chain rule

OpenStudy (anonymous):

dy/dx =2u ,du/dx==4 and dy/dx= (dy/du)(du/dx)=2u(4)= 8u

OpenStudy (anonymous):

did you understand my solution purple?

OpenStudy (anonymous):

we didnt learn to do it like that we do f'(x) = f'(g(x))g'(x)

OpenStudy (anonymous):

ok that is the same...lol.isnt that dy/dx is very simple?

OpenStudy (anonymous):

nope its 8(4x+7)

OpenStudy (anonymous):

@ least its how i learned how to do it

OpenStudy (anonymous):

dy d(u^2) du d(4x+7) ---- = -----=2u , --- = ------ 4 du du dx dx

OpenStudy (anonymous):

du/dx =4

OpenStudy (anonymous):

now dy dy du ---= --- ----= 2u(4)= 8u ans.. dx du dx

OpenStudy (anonymous):

is it much better now purple?

OpenStudy (anonymous):

no its the same...i'll just do that one the way i learned and see what my lectuer says

OpenStudy (anonymous):

i didnt see that bendt did it up there..lol bec my pc keep hanging up..lol

OpenStudy (anonymous):

but you problem reqiured dy/dx, du/dx and dy/du instead of using f(x)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok if you want to use functions dy/du=f '(u)=2u du/dx=f '(x)=4 dy/dx=f '(u) f '(x)= 2u(4)=8u

OpenStudy (anonymous):

hope you get it now?..lol

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you so much